# A Short Proof of the Poincaré Conjecture

### v3

@@ Abstract.
A short, fairly self-contained proof is given of the Poincaré Conjecture. There is a possible error on Page 8. Hopefully it may just be a gap that can be filled.
@@ 1. Introduction
My understanding of simple closed curves on tetrahedrons has benefitted from correspondence with Sam Shepherd and with Andy Bartholomew.
@@ 3. The Proof
We see that $W_{0}$ is the set of vertices of a normal $2$-sphere $S_{0}=\delta M_{0}$. In $S_{1}$ the pair $u_{1},v_{1}$ will be the ends of a returning arc in at least one $2$-simplex. In another $2$-simplex $u_{1},v_{1}$ will be joined by lines in both $S_{1}$ and $S$ to the vertices of an edge in $S_{0}$. Having such a situation in a $2$-simplex is the only way that removing $u_{1},v_{1}$ will give a normal pattern.
$u_{1}$$v_{1}$xxyywwzz
@@
Figure 9 shows the situation for $u_{1},v_{1}$ which lie in the $1$-simplex $\gamma_{1}$. In $S_{1}$, $u_{1},v_{1}$ will be the ends of a retuning arc in one of the $2$-simplexes containing $\gamma_{1}$. This cannot be the case in every $2$-simplex containing $\gamma_{1}$ as it would mean that $S_{1}$ would not be connected. In at least one $2$-simplex containing $\gamma_{1}$ as a face, $u_{1},v_{1}$ will be joined to the vertices $w,z$ of an edge of $S_{0}$, shown dashed red. This is the only way that removal of the pair will result in the normal $2$-sphere $S_{0}$. In $P$, $w,z$ will be joined to points $x,y$ by lines, shown blue in Figure 9. Let $\nu_{1}$ be the permutation of $W$, which restricts to $(v_{1},x)(u_{1},y)$ for $W\cap\gamma_{1}$ and is the identity map on all the other $1$-simplexes. Under the corresponding homotopy the two lines become uncrossed if they were crossed before. Note that the points $x,y$ are adjacent and stay adjacent in subsequent homotopies because the homotopy takes place in $M_{0}$ and so the new lines of the homotopy do not intersect the lines of $S_{0}$. The pattern for $S_{0}$ in each $3$-simplex is as in Figure 1 and so the homotopy $\nu_{1}$ must take place in one of the regions which is not bounded by two parallel lines. This is illustrated in Figure 10, which indicates what can happen in the faces of a $3$-simplex. A $4$-track has been replaced by two $3$-tracks. If instead of joining the ordered pair $u_{1},v_{1}$ to the vertices of the second edge, we joined $v_{1},u_{1}$ to those vertices we would get a $6$-strack. To get a subgraph of $P$ one must join up the points to give two $3$-tracks.
If a $2$-simplex $\sigma$ contains $\gamma_{1}$ as one of its faces, but has no intersection with $S_{0}$, then the points will be joined by a returning arc in $\sigma$.
@@
We have defined $\nu_{1}$. Let $\mu_{1}=\nu_{1}$. We now define $\nu_{i}$ and $\mu_{i}$ for $i=2,\dots,k$ to be permutations of $W$, where $\nu_{i}$ restricts to a permutation on the points of $W\cap\gamma_{i}$ and is the identity on the other points of $W$, and so that the homotopy associated with $\mu_{i}=\nu_{i}\mu_{i-1}$ moves the points of $W_{i}$ to the positions they should have in $P$. In at least one of the $2$-simplexes containing $\gamma_{i}$ as a face there are lines in $S$ joining $u_{j}$ to $u_{i}$ for some $j<i$ or there is a situation as in Figure 9, in which $u_{i},v_{i}$ are joined to the vertices of an edge of $S_{j}$ for $j<i$. This is because there cannot be a returning arc joining $u_{i}$ and $v_{i}$ in every $2$-simplex containing $\gamma_{i}$, as if there were then $S_{i}$ would not be connected.
If there are lines in $\mu_{i-1}S$ joining $u_{i}$ to $u_{j}$ and $v_{i}$ to $v_{j}$ it will be as in Figure 11(a) or there will be three labelled pairs joined by lines as in Figure 11(b). This follows because $u_{i},v_{i}$ is a removable pair in $S_{i}$ and so it will have a returning arc joining $u_{i}$ and $v_{i}$ and this returning arc was created by one or more homotopy $f_{j}$ for $j<i$.
In Figure 11 (a) we see what happens in a $2$-simplex $\sigma_{i}$ containing $\gamma_{i}$, if the lines joining $u_{i}$ and $v_{i}$ come from a single pair $u_{j},v_{j}$. I think there might be a problem with this Figure. We need both $x,y$ to be in $\gamma_{i}$. It might be the case that both points are in the third edge of the $2$-simplex.
In Figure 11(b) the lines to $u_{i},v_{i}$ come from two pairs. The lines in blue are those of $\mu_{i-1}S\cap\sigma_{i}$. The lines in red are the ones we move them to in $\mu_{i}S$. The bottom edge is the $1$-simplex $\gamma_{i}$ containing the pair $u_{i},v_{i}$ which have labels $i,i$ in the labelling for $S$. In at least one of the $2$-simplexes containing $\gamma$ there will be a $1$-face different from $\gamma$ containing an adjacent pair with labels $l,l$ for $l<i$, which are joined by non-intersecting lines in $P$ to the pair $x,y$ in $\gamma_{i}$, or the situation will be as in Figure 11(b) so that there are pairs with labels less than $i$ in each of the other two faces. The induction step is to apply the permutation $\nu_{i}=(v_{i},x)(u_{i},y)$ to the set $\gamma_{i}\cap W$. Of course if $u_{i}=y$ we apply the transposition $(v_{i},x)$. Again with each transposition we also transpose all attaching lines in all $2$-simplexes containing $\gamma$. This change will not disturb any of the vertices of $W_{i-1}$ or the edges joining them. Note that $x,y$ will have labels bigger than $i$ as if, say, $x\in W_{i-1}$, then in $S_{i-1}$ there would be a point in the intersection of a $1$-simplex with $\mu_{i-1}S$ joined to both $u_{i}$ and $y$ and this does not happen in a spattern.
@@
We now know that $\mu S$ is a patterned surface. All the $2$-pieces, apart from the exceptional one, intersect each $1$-simplex at most once and so are $3$-sided or $4$-sided. The $8$-track shown in Figure 6(a) is the only possibility for the exceptional $2$-piece. Thus $S$ has become an almost normal $2$-sphere and we have a proof of the Poincaré Conjecture. In the homotopy $F$, the first thick sphere $S$ will also be the last thick sphere. After applying $\mu$, then for $j>k$ each step of the homotopy becomes an isotopy in which a returning arc joining adjacent points is removed. There is now a new labelling of $W$ in which every point receives a label $j$, where $n+1\leq j\leq n$. The pair of points labelled $j$ will be a removable pair in $f_{j}(S^{2})$.

### v4

@@ Abstract.
A short, fairly self-contained proof is given of the Poincaré Conjecture. In the previous version there was an error on Page 8. This gap has now been filled.
@@ 1. Introduction
My understanding of simple closed curves on tetrahedrons has benefitted from correspondence with Sam Shepherd and with Andrew Bartholomew.
@@ 3. The Proof
As a spattern is determined by its intersections with $2$-simplexes, we will consider what happens to a single $2$-simplex in the homotopy sequence. Initially the $2$-simplex will intersect $S_{0}$ as in Figure 9(a) or Figure 9(b). The unshaded parts will be the intersection with $M_{0}$. Note that since $M_{0}$ contains no vertices, each vertex is contained in a shaded region. The intersection with $S$ will consist of the intersection with $S_{0}$ together with straight lines joining edges in the unshaded regions. The extra intersection points with an edge are paired - each pair lying in an unshaded component.
If our sequence of homotopies was a sequence of isotopies, then each isotopy would result in an increase in shaded area. Thus we could go from Figure 9(a) to Figure 9(b), since removing the pair $u_{i},v_{i}$ from 9(b) gives 9(a). For each $2$-simplex, such a move can occur just once as the central region, which is initially unshaded becomes shaded. If this region is initially shaded then, obviously, no such move can occur. The other move that can occur is adding a shaded region to an unshaded region as shown in Figure 10.
It will be shown that our sequence of homotopies can be converted to a sequence of isotiopies by using the homotopies corresponding to permutations of the points of $W$ on a $1$-simplex.
Let $W_{0}$ ibe the set of vertices of the normal $2$-sphere $S_{0}=\delta M_{0}$. In $S_{1}$ the pair $u_{1},v_{1}$ will be the ends of a returning arc in at least one $2$-simplex. In another $2$-simplex $u_{1},v_{1}$ will be joined by lines in both $S_{1}$ and $S$ to the vertices of an edge in $S_{0}$. Having such a situation in a $2$-simplex is the only way that removing $u_{1},v_{1}$ will give a normal pattern. Thus we are in the situation of Figure 9, with $i=1$,rather than Figure 10. In the sequence of homotopies the lines joining $u_{1},v_{1}$ to the ends $w,z$ of an edge may cross. They can be uncrossed by transposing $u_{1}$ and $v_{1}$. We need $\nu_{1}$ to do more than this. We permute the points of $W$ that lie in the interval $[p,q]$ so that there are no points in the open interval $(u_{1},v_{1})$ and there are no lines that cross the lines from either $u_{1}$ or $v_{1}$. This will happen if the points in $W\cap[p,q]$ are permuted so that the ones joined to a point on the bottom edge are in $q,u_{1}]$ and the the points in $p,v_{1}]$ are joined to points in the right hand edge. The lines $u_{1}w$ and $v_{1}z$ will now be lines in $P$.
[…]
@@
We have defined $\nu_{1}$. Let $\mu_{1}=\nu_{1}$. We now define $\nu_{i}$ and $\mu_{i}$ for $i=2,\dots,k$ to be permutations of $W$, where $\nu_{i}$ restricts to a permutation on the points of $W\cap\gamma_{i}$ and is the identity on the other points of $W$, and so that the homotopy associated with $\mu_{i}=\nu_{i}\mu_{i-1}$ moves the points $u_{i},v_{i}$ to the positions they should have in $P$. In at least one of the $2$-simplexes containing $\gamma_{i}$ as a face there are lines in $S$ joining $u_{j}$ to $u_{i}$ for some $j<i$ or there is a situation as in Figure 9, in which $u_{i},v_{i}$ are joined to the vertices of an edge of $S_{j}$ for $j<i$. This is because there cannot be a returning arc joining $u_{i}$ and $v_{i}$ in every $2$-simplex containing $\gamma_{i}$, as if there were then $S_{i}$ would not be connected.
In Figure 10 (a) we see what happens in a $2$-simplex $\sigma_{i}$ containing $\gamma_{i}$, if the lines joining $u_{i}$ and $v_{i}$ come from the pair $u_{j},v_{j}$.
We permute the points of $W$ that lie in the interval $[r,s]$ so that there are no points in the smaller interval $u_{i},v_{i}]$ and there are no lines that cross the lines from either $u_{i}$ or $v_{i}$. This will happen if the points in $W\cap[r,s]$ are permuted so that the ones joined to a point in $x,u_{j}]$ are in $r,u_{i}]$ and the the points in $s,v_{i}]$ are joined to points in $v_{j},y]$. The lines $u_{i}u_{j}$ and $v_{i}v_{j}$ will now be lines in $P$. A similar argument can be used if both pairs are in the central region.
@@
We now know that $\mu S$ is a patterned surface. All the $2$-pieces, apart from the exceptional one, intersect each $1$-simplex at most once and so are $3$-sided or $4$-sided. The $8$-track shown in Figure 7 is the only possibility for the exceptional $2$-piece. Thus $S$ has become an almost normal $2$-sphere and we have a proof of the Poincaré Conjecture. In the homotopy $F$, the first thick sphere $S$ will also be the last thick sphere. After applying $\mu$, then for $j>k$ each step of the homotopy becomes an isotopy in which a pair of adjacent poiints joined by a returning arc is removed. There is now a new labelling of $W$ in which every point receives a label $j$, where $n+1\leq j\leq n$. The pair of points labelled $j$ will be a removable pair in $f_{j}(S^{2})$.